OpenStudy (anonymous):

What is the first step in solving the equation one-tenth times x minus three-tenths equals seven-tenths? Explain your reasoning.

OpenStudy (anonymous):

You move the -3/10 to the other side of the equation in order to simplify your equation.

OpenStudy (anonymous):

(1/10)x - (3/10) = (7/10) (1/10)x - (3/10) + (3/10) = (7/10) + (3/10) must do same action to both sides of the equation, then rewrite: (1/10)x = (7/10) + (3/10)